A rectangular shaped mansion where the length was 12 more than the square of the width. If the perimeter of the mansion was 168 meters, what is the measure of its area?
a. 1440 sq. meters c. 1420 sq. meters
b. 1430 sq. meters d. 1410 sq. meters

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

A rectangular shaped mansion where the length was 12 more than the square of the width. If the perimeter of the mansion was 168 meters, what is the measure of its area?
-------------------
P = 2W + 2L = 168
W + L = 84
---
L = W^2 + 12
----
W^2 + 12 + W = 84
W^2 + W - 72 = 0
(W + 9)*(W- 8) = 0
W = 8
L = 76
Area = 608 sq meters

Answer by ikleyn(52777)

.

                        D U P L I C A T E


I just solved it yesterday under this link

https://www.algebra.com/algebra/homework/word/unit_conversion/Unit_Conversion_Word_Problems.faq.question.1187413.html

https://www.algebra.com/algebra/homework/word/unit_conversion/Unit_Conversion_Word_Problems.faq.question.1187413.html